OpenStudy (anonymous):

Simplify by rationalizing each denominator.. Square root of 11 over 5 times the square root of 132

OpenStudy (anonymous):

\[\frac{\sqrt{11}}{5\sqrt{132}}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok since \(132=2^2\times 3\times 11\) if you multiply top and bottom by \(\sqrt{3\times 11}\) aka \(\sqrt{33}\) you will have a perfect square in the denominator

OpenStudy (anonymous):

\[\frac{\sqrt{11}}{5\sqrt{132}}\times\frac{\sqrt{33}}{{\sqrt{33}}}\]is a start

OpenStudy (anonymous):

then the denominator will be \[5\sqrt{2^2\times 3^2\times 11^2}=5\times 2\times 3\times 11=330\]

OpenStudy (anonymous):

and the numerator will be \(11\sqrt3\) then you can cancel an 11 top and bottom

OpenStudy (anonymous):

Im confused.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

if you had \[\frac{2}{\sqrt3}\] and you wanted to rationalize the denominator you would multiply top and bottom by \(\sqrt3\) and get \[\frac{2}{\sqrt3}\times \frac{\sqrt3}{\sqrt3}=\frac{2\sqrt3}{3}\]

OpenStudy (anonymous):

is that confusing or is that clear?

OpenStudy (anonymous):

Well i understand how to do that one, but its the 5 thats throwing me off .-.

OpenStudy (anonymous):

ignore the 5, just leave it there

OpenStudy (anonymous):

suppose you had \[\frac{7}{\sqrt{12}}\] and you wanted to rationalize the denominator what you would multiply top and bottom by ?

OpenStudy (anonymous):

The square root of 12

OpenStudy (anonymous):

you could but you do not need to, that is the point here \[12=4\times 3\] and \(\sqrt{4}=2\) so you could just multiply by \(\sqrt3\) top and bottom and get \[\frac{7}{\sqrt{12}}\times \frac{\sqrt3}{\sqrt3}=\frac{7\sqrt3}{\sqrt{36}}=\frac{7\sqrt3}{6}\]

OpenStudy (anonymous):

So you multiply the top & bottom by the multiple of the bottom that doesnt have a square root?

OpenStudy (anonymous):

exactly !

OpenStudy (anonymous):

i mean you could do it the other way as well, but then you would have a bunch to cancel at the end

OpenStudy (anonymous):

that is why the very first thing i wrote was \[132=2^2\times 3\times 11\] so show that you only have to multiply by \(\sqrt{11\times 3}\)

OpenStudy (anonymous):

But in the original question the square root of 11 over 5 times the share root of 132, you multiply the top & bottom but the square root of 33 which gives you 363 on top & 5 times the square root of 4356, thats a big number is that wrong what next? .-.

OpenStudy (anonymous):

if you leave it in factored form it is much easier

OpenStudy (anonymous):

\[\frac{\sqrt{11}}{5\sqrt{2^2\times3\times 11}}\times\frac{\sqrt{33}}{{\sqrt{3\times 11}}}\]

OpenStudy (anonymous):

then you see right away that the denominator is \[5\times 2\times 3\times 11\]

OpenStudy (anonymous):

and the numerator is \[\sqrt{11\times 11\times 3}=11\sqrt3\]

OpenStudy (anonymous):

Im never going to get this! 😥

OpenStudy (anonymous):

it is not that bad really just don't multiply out the whole point is to get a perfect square in the denominator i have no idea what \[2^2\times3^2\times 11^2\]is but i am certain that its square root is \(2\times 3\times 11\)

OpenStudy (anonymous):

But why 2 if its 2^2

OpenStudy (anonymous):

what i meant was \[\sqrt{2^2\times 3^2\times 11^2}\]is \[2\times 3\times 11\] clear or no?

OpenStudy (anonymous):

Wait! I think i have it!

OpenStudy (anonymous):

If the denominator is 5 times the square root of 2^2 x 3 x11 and your multiplying by 3 x 11 it turns into 5 x 2^2 x 3^3 x 11^11 then the square roots of that are 5x2x3x11 ?

OpenStudy (anonymous):

yes yes yes and yes

OpenStudy (anonymous):

But then whats the answer i dont know where to go from that

OpenStudy (anonymous):

ok we write it second to last step is \[\frac{\sqrt{11}\sqrt{3\times11}}{5\times 2\times 3\times 11}\]

OpenStudy (anonymous):

then the numerator is \(11\sqrt3\) so we get \[\frac{11\sqrt3}{5\times 2\times 3\times 11}=\frac{\sqrt3}{5\times 2\times 3}\]

OpenStudy (anonymous):

notice the 11 canceled if we had been paying attention we could have cancelled it first , but no matter